OpenStudy (jakeybabs):

25=7+3x-12 what is x

OpenStudy (anonymous):

25 = 7+3x-12 25 = 3x - 5 30 = 3x 10 = x therefore, x = 10

OpenStudy (jakeybabs):

thanks! had to make sure i was doing this write lol

OpenStudy (jakeybabs):

oops, i meant right*

OpenStudy (anonymous):

no worries, its always good to check :)